OpenStudy (anonymous):

Compute the following limits using L'Hopital's Rule if appropriate. \[\lim_{x \rightarrow \infty} \frac{ \ln(x^5-9) }{ \ln(x) }\] and \[\lim_{x \rightarrow \infty}\frac{ e ^{9x} }{ e ^{10x}-e ^{-10x} }\]

OpenStudy (anonymous):

How do you figure out if the limit is infinity/infinity?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

plug in the value of x, \(\ln \infty = \infty \\ e^\infty = \infty \)

hartnn (hartnn):

so both the limits are indeed of the form infinity/infinity, L'Hopital Rule can be used.

OpenStudy (anonymous):

(ln(x^(5)-9))' = 5x^4/x^5-9 (ln(x))'=1/x

hartnn (hartnn):

yes. now check whether this is on the form of infinity/infinity or 0/0 note: 1/ infinity = 0

OpenStudy (anonymous):

So the 5x^4/(x^-9) is going to infinity and 1/x is 0

OpenStudy (anonymous):

So the final answer is 0??

hartnn (hartnn):

the numerator is 5x^4/(x^5-9) is equivalent to 5x^4/x^5 for large values of x, so, 5/x or 5/infinity = 0 so, the limit is indeed of the form 0/0 makes sense?

OpenStudy (anonymous):

Okay.. Kind of get it